A man moves a mass of 95kg with a velocity of 15.3m/s. After sometime, he suddenly bumped into a big box and released the energy of 2013 J as work. What is now the new velocity of the mass after the event


1
Expert's answer
2021-09-06T11:33:12-0400

The energy-work theorem gives

"\frac{mv_2^2}{2}-\frac{mv_1^2}{2}=W"

Hence, the new velocity of the mass after the event

"v_2=\sqrt{v_1^2+2W/m}\\=\sqrt{(15.3\:\rm m/s)^2+2*(-2013\:\rm J)/95\:\rm kg}\\=13.8\:\rm m/s"